1. There are stations having the maximum frequency deviation of 75 kHz, and the maximum audio frequency for the modulation is 15 kHz , find its deviation ratio.
  2. Determine the peak frequency deviation for a FM modulator with a deviation sensitivity of 12 kHz/V  and a peak modulating signal of 9V.
  3. If the maximum frequency deviation of the carrier is 10 kHz and the maximum modulating frequency is 6.7 kHz, find the modulating index.
  4. A PM signal has a maximum allowed deviation of 22 kHz  and an actual deviation of 5 kHz . What is its percent modulation?
1
Expert's answer
2021-10-25T15:29:40-0400

Deviation ratio

= "\frac{maximum allowable deviation}{maximum modulating frequency}"

="\frac {75}{15}"

=5


Frequency deviation =Am kf

Where Am = modulating frequency

kf= frequency sensitivity

Frequency deviation = 12×9

108kHz


Modulating index="\frac {frequency deviation}{modulating frequency}"

= "\frac {10}{6.7}"


=1.4925



Percentage modulation= "\frac {actual frequency deviation}{allowable deviation}"


="\frac {5}{22}×100\%"


=22.73%